Percentage:
I have 49 apples.
38 of the apples are red.
What is the percentage of red apples?

Answer by tutorcecilia(2144)

Part/Whole = Percent/100 [Use the formula for percentages]
38/49 = percent/100 [Plug-in the values]
38/49=x/100 [solve for x]
(38)(100)=(49)(x)
3800/49=77.55%
.
Check by plugging in 77.55% back into the formula for percentages:
38/49=77.55/100
.7755=.7755 [Checks out]